A force of 20N act on an object of an angle 30degrees to the inclined plane. If the object moves 8metres up the plane

(a) What is the effective force that moves the object up the plane

(b) what is the work done to move the object up the plane


1
Expert's answer
2021-12-27T08:02:59-0500

(a) The effective force is


"F=20\cos(30°-\theta),"

where theta is the angle of the plane above the horizontal.

(b) The work is


"W=Fd=20\cos(30°-\theta)·8=160\cos(30°-\theta)."